Calculate the price elasticity of demand if the price changed from $13 to $12 and from $2 to $1


1
Expert's answer
2021-01-07T08:34:22-0500


Price elasticity in demand is defined as change in demand relative to the prices and can be calculated in mathematical terms to determine their impact in the economy.

It is calculated by dividing percentage quantity demanded and the percentage change in overall prices.

The calculation is as follows;


"(12-13)\/(12+13)=-1\/25"

(1-2)/(1+2)=-1/3

=3/25

=0.12